OpenStudy (dmezzullo):

Solve: 4(x + 1) = −4 0 −2 2 −1.25

OpenStudy (anonymous):

Do u know the distributive property?

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

4(x+1)=-4 4x+4=-4 -4 -4 4x=-8 4x/4=-8/4 x=-2 <-----answer

OpenStudy (anonymous):

yes ur right

OpenStudy (hba):

@labayse Please Do not give direct answers,it is against the CoC - http://openstudy.com/code-of-conduct

OpenStudy (dmezzullo):

i got it nvm